If a new law decreases the highest marginal tax rate for all companies from 35% to 21% for next fiscal year. What would be the forecasted WACC? Market Risk premium of 4.4%. 1 year treasury yield curve of 2.6%. Beta of .52. Please assume the proportion of debt and equity, and also the cost of debt. Any other information missing must be assumed. Please show formula used and work. What is the new WACC equal to? WACC=?
Answer #1

Assume Weight of debt = + Equity)   0.5
Cost of debt assume   3%
Weight of equity = 1-0.5=   0.5
  
Cost of Equity = Risk free rate +(Beta*(Market RISK premium.)  
2.6% +(0.52*4.4%)=   4.89%
New tax rate=   21%
after Cost of debt = Pretax cost * (1-tax rate)  
3%*(1-21%)=   2.37%
  
WACC = (weight of debt * cost of debt) + (weight of equity * cost of equity)  
(0.5*2.37%) + (0.5 * 4.89%)  
3.63%  
  
So, new WACC is 3.63%